Instrument Log

 

SurTHRef C1 Lamont, OK

 

Date

Technician

Comments

9/22/14

JEM

SurTHRef instrument decommissioned as per BCR 02046.

8/6/14

JEM

The RH reading from probe R3 has dropped out. Checked all connections for R3 and they were all OK. R/R *R3 @1335 GMT. Edited the datalogger program to reflect the change and verified all readings were OK. RTS @1445 GMT. *Old - 92781 *New - 92783

4/21/14

JEM

6-month calibration needs to be completed. Performed 6-month calibration using the Vaisala HMP75 handheld probe B2060003 with a calibration date of 01/07/2014. All sensors were within specifications on temperature and humidity and did not require replacement. RTS @1915 GMT.

4/14/14

CAM

DS View page indicated no data had been collected from SurTHRef for 13 hours. T/S found LoggerNet was closed. Opened LoggerNet at 1302 GMT. All missing data successfully recovered on the following collection at 1402 GMT.

3/31/14

PAD

SurTHRef caps need to be cleaned. Removed caps at 1743 GMT and cleaned. Installed at 1758 GMT. Temperature and RH values returned to normal by 1813 GMT.

1/3/14

KT

The SurTHRef was not reporting any data. At 1350 GMT troubleshooting found that one of the batteries located inside the CR2300 was reading -2.2 VDC and needed to be changed out. After changing out the battery the instrument was powered on. RTS at 1450 GMT.

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm. Power was restored at 1920 GMT. Instrument RTS status O.K and data collections resumed once SDS operations were restored.

1/28/13

PAD

SurthRef caps need to be cleaned. Removed caps from all Temp/Hum probes at 1418 GMT and cleaned. Installed caps at 1445 GMT. RTS at 1448 GMT.

10/22/12

KT

There was no communications with the instrument. At 1325 GMT power cycled the NL100. At 1530 GMT communications were established with instrument. RTS at 1340 GMT.

8/22/12

PAD

SurTHRef sensor caps need to be cleaned. At 1350 GMT, removed sensor caps and washed with water. Removed excess water with canned air and installed caps. RTS at 1410 GMT.

7/2/2012

JEM

R3 needs to be replaced for calibration. R/R R3 @2035 GMT. Edited and reloaded the datalogger program to reflect the change and RTS @2045 GMT. *Old 92783 *New - 92781

4/23/12

JEM

R1 needs to be replaced and sent in for calibration. R/R *R1 @1745 GMT. Updated program @1751 GMT and verified readings. RTS @1802 GMT. *Old - 92781 *New 92784

2/16/12

JEM

Since V2 was installed there have been multiple times where RH readings have dropped offscale. It appears that during high humidity conditions the RH sensor is going as high as 110% and then dropping offscale (-6999). Once the RH starts to drop to lower levels, the sensor will return to a reasonable reading. Sensor needs to be returned for repair and possibly calibration. Replaced *V2 @1930 GMT. Edited the serial number in the program and loaded it to the logger @1937 GMT. Readings appeared reasonable. RTs @1940 GMT.

*Old - Y31100014 *New - S4930005

1/24/12

JEM

R2 is due to be replaced for calibration. R/R *R2 @2137 GMT. Readings were observed reasonable after replacement. Updated the program to reflect the new serial number @2151 GMT. RTS @2155 GMT.

*Old 92784 *New 92782

1/6/12

JEM

V1 and V3 need to be changed out for annual calibration. R/R *V1 and **V3 beginning at 1850 GMT. After replacing the sensors, updated the program to reflect the new serial numbers. Loaded the new program at 1912 GMT. All readings appeared reasonable after replacement. RTS @1920 GMT. *Old - V2730033 *New - Y3110012 **Old - Y3110013 **New - Y3110015

12/20/2011 

JEM

Sensor V2 RH values are going off scale to -6999.  Checked and re-seated all connections.  The reading is still -6999.  Removed the sensor cap and unplugged the RH sensor from the end of the probe.  Slightly spread apart the sensor legs then plugged it back into the probe.  RH readings have now returned.  RTS @1525 GMT.

12/12/201

JEM

All sensors are due for calibration.  Due to limitations on spares sensors will be rotated out as spares become available.  Replaced *V2 @2100 GMT. The new serial number is Y3110014. It is not due for calibration until 08/2012. Updated the program to reflect the new serial number @2112 GMT.  RTS @2115 GMT. *Old - Y3110012 *New - Y3110014

12/2/11

PAD

Removed and cleaned Temp/RH sensor caps at 1436 GMT. Replaced caps at 1456 GMT.

11/10/2011

JEM

Mentor requests that sensor *V1 be replaced and sent in for calibration due to consistently high RH values. Replaced the sensor between 1820 GMT and 1852 GMT. Verified the readings were reasonable. Updated the program to change the serial number for V1 from 0015 to 0033. Loaded the new program @1912 GMT. RTS @1915 GMT.

*Old- Y3110015 *New V2730033

9/20/11

PAD

Last SDS collection was at 1400 GMT. Computer is unable to connect with datalogger. Troubleshooting with computer department found datalogger and NL100 convertor had correct voltage. Screen is blank on 23X datalogger. Power cycled datalogger and NL100 and datalogger display is now working. Checked Loggernet status for connections and it alternates between off line, pending, and on line. Brought up Surthref display screen and computer is now connected with datalogger. Water had pooled in bottom of datalogger box. Removed water and repositioned conduit suspected of funneling water into box. RTS at 1745 GMT.

8/24/11

PAD

Cleaned Temp/Hum probe caps at 1352 GMT.

7/11/11

PAD

Unable to connect with datalogger. Used task manager to end loggernet program. Restarted program and now able to connect with datalogger. RTS at 1233 GMT.

3/25/11

MK

Mentor has indicated that V1 is biasing high during saturated or near saturated conditions. R/R *V1 at 1600 GMT. Edlogged program and reloaded into CR23 at 1618 GMT. *Old Y3110014 New- Y3110015.

1/5/11

PAD

Removed caps from Temp/ RH probes and cleaned filters at 1457 GMT.

11/15/10

Mentor

SDS has not collected any data from SurTHRef since 2202 GMT on 11/10. Check of computer found data is being generated and displayed but all new data is being appended to a single file in C:\SurTHRefData. The file is not being split and moved to ftphome directory where it can be collected. Loggernet appears to be corrupted as it shows up as running in the task manager, but the GUI cannot be seen to check settings. Restarting program and computer has no effect. Mentor remotely removed Loggernet 3.4.1 at 2045 GMT and reinstalled the same version. Loggernet then functioned properly. RTS at 2115 GMT.

4/19/10

PAD

Surthref computer is experiencing intermediate connection with datalogger. Troubleshooting found issue with Allied Telesyne media converters located at computer and datalogger. R/R media converters. Communication is now restored. RTS at 1905 GMT.

3/2/2010

MK

Mentor has indicated that R3 is biasing high during saturated or near saturated conditions. R/R *R3 at 2100 GMT. Edlogged program and reloaded into CR23 at 2120 GMT. *Old 92782 *New- 92783

2/19/10

PAD

During PM rounds, observed one of the fans that draws air across the Temp/Hum probes in the Stevenson screen has failed. R/R fan. RTS at 2037 GMT.

12/15/09

PAD

Last SDS collection was on 12/14/09 at 1902 GMT. Computer will not connect with datalogger. Power cycled NL100 network interface link and comms were restored. RTS at 1543 GMT.

12/2/09

PAD

Last SDS collection was on 11/30/09 at 0202 GMT. Computer is unable to connect with datalogger. Rebooted computer but still unable to connect with datalogger. Check of CR 23X datalogger shows display is blank. After disconnecting DL power supply, computer connected and stored data was transferred. No problem found with power supply. Reconnected power supply to DL. Will continue to monitor.

11/3/09

MK

Mentor has indicated that R1 is biasing high during saturated or near saturated conditions. R/R *R1 at 1930 GMT. Edlogged program and reloaded into CR23 at 2009 GMT. *Old 92783 New- 92781

9/21/09

MK

A few of the sensors seem to be reading high, mentor has suggested replacing and cleaning the filter caps. Replaced all three Viasala Temp/hum caps and cleaned the three Rotronic Temp/hum caps at 1610 GMT. RTS at 1625 GMT.

4/20/09

MK

Mentor has indicated that R2 is biasing high during saturated or near saturated conditions. R/R *R2 at 1950 GMT. Edlogged program and reloaded into CR23 at 2005 GMT. *Old 92781 New - 92784

10/20/08

MK

The V3 T/H sensor frequently measures over 100% humidity on rainy days and even drops out completely to a negative value. The sensor needs to be replaced. R/R *V3 sensor, edlogged program and re-loaded onto SurTHRef datalogger. *Old - Y3110015 *New - Y3110013

8/7/08

MK

Two SURTHREF T/H probes are scheduled for re-calibration. Removed *V3 S/N Y3110013 and installed S/N Y3110015. Removed **R1 S/N 92784 and installed S/N 92783. Edlogged SurTHRef program, loaded program and restarted system at 1937. RTS at 1945 GMT.

6/24/08

MK

Two T/H sensors have returned from factory calibration and need to be swaped out with two currently installed sensors so they can be sent in for re-calibration. Removed *R3 Rotronic MP100H sensor S/N 92783, installed S/N 92782 from 1525-1535 GMT. Removed *V1 Vaisala HMP45D sensor S/N Y3110015 , installed S/N Y3110014 from 1612-1640 GMT. Edlogged SurthRef program and reloaded into CR23X datalogger and restarted at 1648 GMT. RTS at 1655 GMT.

3/25/08

MK

The Mentor has requested that we swap out the T/H sensors and return them to vendors for calibrations. Removed the *V2 sensor S/N Y3110014, and installed S/N Y3110012 in it's place. Removed the *R2 sensor S/N 92782 and installed S/N 92781 in it's place. Edlogged program for proper S/N's installed, loaded program into CR23X and returned to service. RTS at 1900 GMT.

3/18/08

CAM

Removed Rotronics sensors from their mounts momentarily to record serial numbers between 2030 and 2035 GMT.

11/6/07

TMG

Removed the previously installed time protocol program and installed Domain Time II on BBSS computer at 2056 GMT and on the SurTHRef computer at 2030 GMT.

10/26/07

MK

The V3 T/H sensor is reporting faulty humidity readings. Removed probe V3 from SurTHRef housing, cleaned removable sensor connections, still reporting 0.0 as the current humidity. Removed HMP45D sensor and replaced. New S/N is Y3110013, Old S/N is Y3110012, edlogged program for replacement of sensor and uploaded. All readings have returned to normal. RTS at 1805 GMT.

5/22/07

MK

The SurTHRef is not transferring data files at this time. Monitored CR23X all values are updating. Power cycled CR23X at 1855 GMT. At 1902 GMT the instrument was transferring data files. RTS at 1915 GMT.

5/21/07

MK

The SurTHRef Temp/Hum probes are scheduled to be changed out one at a time for factory recalibration. R/R *R1 sensor at 1600 GMT. R/R **V3 sensor at1615 GMT. Edlogged program and loaded on to the CR23X at 1620 GMT. *Old - 92781 *New - 92784 **Old 0013 **New - 0012

3/7/07

MK

The SURTHREF S/N's for V2 and V3 sensors are incorrect in the program and need to be switched. Edlogged program, changed V2 to 14 and V3 to 13. Loaded program into CR23X at 1925 GMT.

8/1/06

MK

SURTHREF V1 sensor is reading a temperature and humidity of 9999.0. Removed V1 sensor from enclosure to inspect removable sensor connector. Corrosion was found all over the pins and plug. Cleaned all corrosion from sensor and tested. Readings were still out of tolerance. R/R HMP45D sensor at 1530 GMT. Edlogged SurTHRef program and loaded into CR23X datalogger. New S/N is Y3110015, old S/N is Y3110012.

5/11/06

MK

The Viasala HMP45D probes installed in the SURTHREF have been corroding due to excessive moisture intruding into the electronics of the sensor. In an attempt to correct this problem RTV has been placed around the inside of the tension relief portion of each probe.

5/2/06

MK

On 5/2/06 Instrument mentor has requested that any SURTHREF sensors calibrated in 2003 should be sent in for recalibration. The HMP45D sensor currently installed in the V1 location on the SURTHREF needs to be changed out and sent in. The SURTHREF monitor is currently displaying 9999 for the V1 and V3 humidity sensors. On 5/2/06 disassembled the V1 sensor, drained all water, cleaned out all corrosion. Reading returned to normal. Disassembled V3 sensor noticed that corrosion has damaged one pin on the removable sensor. On 5/4/06 removed V3 sensor at 1930 GMT, replaced damaged pin on sensor plug and cleaned out all corrosion, reinstalled sensor. Humidity reading returned to normal. R/R *V1 HMP45D sensor for recalibration at 2015 GMT. *Old - Y3110015 *New - Y3110012

12/12/05

MK

The temperature reading from the V1 sensor on the SurTHRef system is reading -2.9 C. The other 5 sensors are reading in the range of 12-13C. Found a loose wire connection on the CR23X, tightened connection. Reading returned to normal at 1930 GMT.

12/1/05

MK

T/RH probe Vaisala #3 is failing on the SurTHRef system. Disassembled HMP45D S/N 13 probe that was installed in V3 terminals to check for corrosion on the removable probe connector. Corrosion was found, removed all corrosion from connector and installed probe S/N13 back into the V3 terminals on the CR23X. All readings returned to normal. RTS at 2025 GMT.

10/18/05

MR

DQPR # 601 requests that the SurTHRef *T/RH probe # 1 be checked and or changed out due to it reading -9999 for the humidity. Checked the connections and they were tight. Located the spare probe and changed this probe out. Humidity and temperature readings came back to normal. RTS at 2045 GMT. *Old - Y3110012 *New - Y3110015

10/13/05

TMG

The SurTHRef has a Dell GX270 computer which is known

to have motherboard failures. Mike Ritsche requested Site Ops R/R the motherboard before it failed. Shutdown computer @ 1741 GMT. R/R motherboard @ 1755 GMT. RTS @ 1813 GMT. Status O.K.

9/ /05

Mentor

SurTHRef installed by instrument mentor.